OpenStudy (141823):

A certain system of a linear and a quadratic equation has two solutions, (2,7) and (5,10). The quadratic equation is y = x^2 - 6x + 15. What is the linear equation?

OpenStudy (141823):

@Owlcoffee

OpenStudy (141823):

@kidrah69

OpenStudy (141823):

@TheSmartOne

OpenStudy (141823):

@NoelGreco

OpenStudy (141823):

@zepdrix

zepdrix (zepdrix):

Hey :)

zepdrix (zepdrix):

So we're given the quadratic, and we need to find the linear function. It will have the form, in slope-intercept form, of:\[\Large\rm y=mx+b\]

zepdrix (zepdrix):

This system has two solutions, meaning, that BOTH equations will pass through the given points. So our linear equation will pass through these two points. We can use our slope formula to find the slope between these two points. \[\Large\rm m=\frac{y_2-y_1}{x_2-x_1}\] Remember how to do that? :)

OpenStudy (141823):

live saver right here, I remember this!

zepdrix (zepdrix):

So what do you get for your slope? :U

OpenStudy (141823):

1?

zepdrix (zepdrix):

Mmm ok good!\[\Large\rm y=mx+b\]\[\Large\rm y=1x+b\]\[\Large\rm y=x+b\]Now we just need to find the starting point, the y-intercept. Plug one of those two given points into the function, and solve for b.

OpenStudy (141823):

b = 5, so the equation is y = x + 5

zepdrix (zepdrix):

Yayyy good job \c:/ Here is what it looks like graphed just in case you were curious. https://www.desmos.com/calculator/mgzzovgxtx

zepdrix (zepdrix):

Woops this link: https://www.desmos.com/calculator/07lq6tizry

zepdrix (zepdrix):

Just a website for graphing :) Kinda useful.

OpenStudy (141823):

Thank you zepdrix! I'll see you around : D